OpenStudy (anonymous):

A triangle has an area of 26 in2, and two of the sides of the triangle have lengths 8 in. and 7 in. Find the angle included by these two sides. (Assume the angle is acute. Round your answer to one decimal place.) ...working...

OpenStudy (anonymous):

26=1/2(8)(7)sin(theta) sin(theta)=(2*26)/(8*7)=52/56 theta= ??? after 52/56 how do i find what theta is? what do i do with sin?

OpenStudy (cwrw238):

sin theta = 52/56 (not theta) use your calculator to find theta [ the required key on your calculator will be sin-1 or arcsin]

OpenStudy (anonymous):

asin(52/56)?

OpenStudy (anonymous):

sweet thanks, 68.2 degrees